# Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security

### v9

@@ 5.3 What can be done?
A persistent feature in the literature is the lack of rigorous analysis of min-entropy (and induced guesswork) under realistic acceptance regions and attacker priors [49, 7, 9, 10, 21, 8, 5, 6]. We urge the community to consider the resilience of proposals to adversarial samples, particularly when using ‘fuzzy’ matching methods such as supervised learning models, and to translate those estimates to concrete costs, such as expected guesses.
@@
User actions such as prescribed shakes or gestures, as well as externally induced signals, can inject variability and modestly raise entropy by several bits [26, 28]. For example, altering the device’s environment to produce distinctive electromagnetic signatures may be detectable by onboard sensors. A key gap is to quantify how much $\Delta H_{\infty}$ such interventions can actually deliver across modalities. Our expectation is that any gains remain modest, but this trade-off between added entropy and practical feasibility remains an under-explored area of research.
@@
Where applications need similarity evidence, such as for device authentication, one should treat that evidence as weak, predictable inputs. The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see [38, 40]). Cryptographic extractors (e.g. Von Neumann and later constructions [56, 57, 58]) or decorrelation techniques (e.g. PCA) may be considered to convert low-quality sensor noise into higher-entropy sources. However, while extractors reduce bias and correlation, they cannot increase intrinsic min-entropy. That is, if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more. Extractors improve the quality of usable randomness when sufficient $H_{\infty}$ already exists; they do not increase the quantity in a way that raises brute-force cost. Future work in the area should focus on how entropy can be meaningfully increased while retaining application utility.
@@ 5.4 Scope and Limitations
While we analyse four large datasets, they do not fully capture contexts all possible contexts. Some work has suggested that performing dedicated movements, such as gestures, can increase the amount of usable entropy from motion sensors in the region of approximately 5–6 bits [28, 26]. Our datasets do not cover such dedicated movements; it is possible that the reported results are an underestimation of entropy for sensors which are deliberately perturbed as an entropy-generating action. Conversely, we do not examine entropy-reducing threats, such as sensor spoofing [10], hardware attacks [44], or cross-device correlation attacks [45]. It is concievable that these threats may reduce the entropy of sensors even further. [^5] By way of example, we could consider a voltage-based fault injection attack that causes nondeterministic errors in the output of individual sensors, affecting the entropy of their output values.
We also estimate unconditional entropies such that all measurements are independent and identically distributed (i.i.d.). This is an optimistic assumption, as sensor data is inherently time-correlated; a sensor’s next reading is often highly dependent on its previous readings. For Shannon entropy and for the conditional Rényi notions standard in cryptography (e.g. min-entropy), the conditional entropy is always less than or equal to the unconditional entropy; that is, $H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see [59] for additional properties). Thus, the unconditional entropy values we calculate in this work act as an upper bound. We may wish to condition $X$ on a sequence of previously observed measurements; however, efficiently quantifying $H_{\alpha}(X_{t}\,|\,X_{<t})$ for high-dimensional sensor streams remains non-trivial. Nevertheless, we posit that this would only serve to decrease entropy, strengthening our core conclusion that mobile sensors are insecure data sources for security applications.
@@ 6 Conclusion
Our findings also challenge the common practice that model evaluation metrics (e.g. accuracy or EER) suffice to demonstrate that sensor-based security applications are, indeed, secure. The vulnerability of mobile sensors to biased distributions and inter-sensor correlations significantly undermines their effectiveness as reliable sources of unpredictability. The effectiveness of countermeasures remains an open and pressing research challenge without compromising security or utility. Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, paving the way for safer and more robust designs.

### v10

@@ 5.3 What can be done?
A persistent feature in the literature is the lack of rigorous analysis of min-entropy (and induced guesswork) under realistic acceptance regions and attacker priors fomichev2019perils , mehrnezhad2015tap , shrestha2014drone , shrestha2018sensor , markantonakis2024using , mayrhofer2009shake , gurulian2017effectiveness , gurulian2018good . We urge the community to consider the resilience of proposals to adversarial samples, particularly when using ‘fuzzy’ matching methods such as supervised learning models, and to translate those estimates to concrete costs, such as expected guesses.
@@
User actions such as prescribed shakes or gestures, as well as externally induced signals, can inject variability and modestly raise entropy by several bits voris2011accelerometers , lv2020analysis . For example, altering the device’s environment to produce distinctive electromagnetic signatures may be detectable by onboard sensors. A key gap is to quantify how much $\Delta H_{\infty}$ such interventions can actually deliver across modalities. Our expectation is that any gains remain modest, but this trade-off between added entropy and practical feasibility remains an under-explored area of research.
@@
Where applications need similarity evidence, such as for device authentication, one should treat that evidence as weak, predictable inputs. The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see suciu2011unpredictable , wallace2016toward ). Cryptographic extractors (e.g. Von Neumann and later constructions von195113 , trevisan2001extractors , barak2003true ) or decorrelation techniques (e.g. PCA) may be considered to convert low-quality sensor noise into higher-entropy sources. However, while extractors reduce bias and correlation, they cannot increase intrinsic min-entropy; if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more. Extractors improve the quality of usable randomness when sufficient $H_{\infty}$ already exists; they do not increase the quantity in a way that raises brute-force cost. Future work in the area should focus on how entropy can be meaningfully increased while retaining application utility.
@@ 5.4 Scope and Limitations
While we analyse four large datasets, they do not fully capture contexts all possible contexts. Some work has suggested that performing dedicated movements, such as gestures, can increase the amount of usable entropy from motion sensors in the region of approximately 5–6 bits lv2020analysis , voris2011accelerometers . Our datasets do not cover such dedicated movements; it is possible that the reported results are an underestimation of entropy for sensors which are deliberately perturbed as an entropy-generating action. Conversely, we do not examine entropy-reducing threats, such as sensor spoofing shrestha2018sensor , hardware attacks shepherd2021physical , or cross-device correlation attacks dautov2019effects . It is concievable that these threats may reduce the entropy of sensors even further. [^5] By way of example, we could consider a voltage-based fault injection attack that causes nondeterministic errors in the output of individual sensors, affecting the entropy of their output values.
We also estimate unconditional entropies such that all measurements are independent and identically distributed (i.i.d.). This is an optimistic assumption, as sensor data is inherently time-correlated; a sensor’s next reading is often highly dependent on its previous readings. For Shannon entropy and for the conditional Rényi notions standard in cryptography (e.g. min-entropy), the conditional entropy is always less than or equal to the unconditional entropy; that is, $H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see iwamoto2013information for additional properties). Thus, the unconditional entropy values we calculate in this work act as an upper bound. We may wish to condition $X$ on a sequence of previously observed measurements; however, efficiently quantifying $H_{\alpha}(X_{t}\,|\,X_{<t})$ for high-dimensional sensor streams remains non-trivial. Nevertheless, we posit that this would only serve to decrease entropy, strengthening our core conclusion that mobile sensors are insecure data sources for security applications.
@@ 6 Conclusion
Our findings also challenge the common practice that model evaluation metrics (e.g. accuracy or EER) suffice to demonstrate that sensor-based security applications are, indeed, secure from a data-centric perspective. The vulnerability of mobile sensors to biased distributions and inter-sensor correlations significantly undermines their effectiveness as reliable sources of unpredictability. The effectiveness of countermeasures remains an open and pressing research challenge without compromising security or utility. Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, paving the way for safer and more robust designs.